OpenStudy (anonymous):

Use the demoivre's theorem to find the answer? (1+i)^8

OpenStudy (amistre64):

this is something like: z = w^8((8)1+ (8)i ) right?

OpenStudy (amistre64):

w = 1 maybe lol

OpenStudy (anonymous):

\[\sqrt(2)\times e^{i\pi/4}\]

OpenStudy (anonymous):

that to power 8

OpenStudy (amistre64):

1,8,28,56,70,56,28,8,1 1 +8i +28i^2 +56i^3 +70i^4 +56i^5 +28i^6 +8i^7 +i^8 using pascals triangle :)